A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Show your work.
Sn: 2 is a factor of n^2 + 7n
pls show step by step answer

Answer by stanbon(75887)

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Show your work.
Sn: 2 is a factor of n^2 + 7n
S1: 2 is a factor of 1^2+7*1 = 1+7 = 8 = 2(4) is even
S2: 2 is a factor of 2^2+7*2 = 2(2+7) is even
S3: 2 is a factor of 3^2+7*3 = 9+21 = 2(15) is even
------------
Cheers,
Stan H.
--------------